# EMOVA[Uncaptioned image]: Empowering Language Models to See, Hear and Speak with Vivid Emotions

### v2

@@ Abstract
GPT-4o, an omni-modal model that enables vocal conversations with diverse emotions and tones, marks a milestone for omni-modal foundation models. However, empowering Large Language Models to perceive and generate images, texts, and speeches end-to-end with publicly available data remains challenging in the open-source community. Existing vision-language models rely on external tools for the speech processing, while speech-language models still suffer from limited or even without vision-understanding abilities. To address this gap, we propose EMOVA (EM-otionally Omni-present Voice Assistant), to enable Large Language Models with end-to-end speech capabilities while maintaining the leading vision-language performance. With a semantic-acoustic disentangled speech tokenizer, we notice surprisingly that omni-modal alignment can further enhance vision-language and speech abilities compared with the corresponding bi-modal aligned counterparts. Moreover, a lightweight style module is proposed for flexible speech style controls (e.g., emotions and pitches). For the first time, EMOVA achieves state-of-the-art performance on both the vision-language and speech benchmarks, and meanwhile, supporting omni-modal spoken dialogue with vivid emotions.
@@ 1 Introduction
OpenAI GPT-4o (OpenAI, 2024), a new milestone for omni-modal foundation models, has rekindled people’s attentions on intelligent assistants that can see (i.e., perceiving fine-grained visual inputs), hear (i.e., understanding vocal instructions) and speak (i.e., generating vocal responses) simultaneously. Most existing Multi-modal Large Language Models (MLLMs) focus on two modalities only, either vision-language (Bai et al., 2023; Li et al., 2024a) or speech-language (Chu et al., 2024; Xie & Wu, 2024), demonstrating severe demands for omni-modal models with visual, language and speech abilities. How to effectively empower Large Language Models (LLMs) to process omni-modal data in an end-to-end manner remains an open question.
@@
Existing omni-modal LLMs (Chen et al., 2024b; Fu et al., 2024b) generally build upon Vision LLMs and integrate the speech modality by adopting a speech encoder like Whisper (Radford et al., 2023), which extracts continuous features from speech, similar to how images are processed, and enables speech understanding. However, these models still rely on external Text-to-Speech (TTS) tools for generating speech responses, limiting their ability to support real-time interactions. AnyGPT (Zhan et al., 2024), instead, opts for a fully discretization manner, which first discretizes all data modalities (i.e., images, texts, and speeches), followed by omni-modal auto-regressive modeling. This enables AnyGPT to handle multiple modalities with a unified end-to-end framework, facilitating real-time interactions with the help of streaming decoding. However, the discrete vision tokenizer adopted by AnyGPT struggles to capture visual details, especially for high-resolution images, making it far behind its continuous counterparts on vision-language benchmarks. Moreover, none of the existing works explore speech style controls (e.g., emotions and pitches) with LLMs. Therefore, our question arises: How to build an end-to-end omni-modal LLM enabling spoken dialogue with vivid emotions while maintaining state-of-the-art vision-language performance?
@@
In this paper, we propose EMOVA (EMotionally Omni-present Voice Assistant), a novel end-to-end omni-modal LLM with state-of-the-art vision-language and speech capabilities while supporting emotional spoken dialogue. Fig. 1 shows an overview of the model framework. A continuous vision encoder captures the fine-grained visual details, while the discrete speech tokenizer and detokenizer enable the end-to-end speech understanding and generation. Specifically, the speech-to-unit (S2U) tokenizer converts the input speech waveforms into discrete speech units as LLM inputs, while the unit-to-speech (U2S) detokenizer reconstructs the speech waveforms from the LLM’s output speech units. To seamlessly integrate the speech modality with LLMs, we meticulously design a semantic-acoustic disentangled speech tokenizer to decouple the semantic contents and acoustic styles of the input speeches (Tao et al., 2024), where 1) semantic content (i.e., what it says) captures the semantic meanings of input speeches, which is finally discretized and aligned with LLMs, while 2) acoustic style (i.e., how it says) captures the diverse speech styles (e.g., emotions and pitches). Utilizing the semantic-acoustic disentanglement of our speech tokenizer, we further introduce a lightweight style module to support spoken dialogue with vivid emotions and pitches. As in Sec. 4.1, this innovative disentanglement design better facilitates the modality alignment between texts and speeches while maintaining flexibility for diverse speech style controllability and personalization.
@@
With the end-to-end omni-modal architecture of EMOVA, we empirically demonstrate that publicly available bi-modal image-text and speech-text data are sufficient for omni-modal alignment, utilizing the text modality as a bridge. This eliminates the need for omni-modal data (i.e., image-text-speech), which is usually scarce. Surprisingly, we find that omni-modal alignment can further improve both vision-language and speech capabilities through joint optimization, even when compared with their bi-modal aligned counterparts. Finally, only a small amount of mixed-modality samples are required to teach the model to respond in the desired format. For the first time, EMOVA achieves state-of-the-art performance on both vision-language and speech benchmarks (see Table 1 for comparisons). The main contributions of this work contain three parts:
Figure 1: Model architecture of EMOVA. The vision encoder extracts continuous visual features, which are projected into the text embedding space as visual tokens, while the input speech is encoded and quantized into discrete units. Given the omni-modal inputs, EMOVA can generate both textual and speech responses with vivid emotional controls. Check Sec. 3 for more architectural details.
@@
    We propose EMOVA, a novel end-to-end omni-modal LLM that can see, hear and speak. We use a continuous vision encoder and a semantic-acoustic disentangled speech tokenizer for seamless omni-modal alignment and diverse speech style controllability.
@@
    We introduce an efficient text-centric omni-modal alignment which can further improve the vision-language and speech capabilities, even compared with the corresponding bi-modal aligned counterparts (i.e., image-text only and speech-text only alignment).
@@
    For the first time, our EMOVA achieve state-of-the-art comparable performance on both the vision-language and speech benchmarks simultaneously, while supporting flexible spoken dialogues with vivid emotions.
Table 1: Comparison among Multi-modal Large Language Models. Our EMOVA is the very first unified Omni-modal Large Language Model capable of emotional spoken dialogue with state-of-the-art vision-language and speech capabilities simultaneously.
@@
Understand Generation Emotional
Vision LLaVA
Intern-VL
Speech Qwen-Audio
Mini-Omni
LLaMA-Omni
Intern-Omni
VITA
Any-GPT
Omni EMOVA (ours)
@@ 2 Related Work
(VLLMs) integrate vision modality into Large Language Models (LLMs) (Touvron et al., 2023; Chen et al., 2023b), enabling advanced understanding and reasoning over visual instructions (Liu et al., 2024b; Bai et al., 2023; Gou et al., 2023; 2024). Recent efforts in VLLMs can be broadly categorized into three directions, including 1) Vision encoders (Oquab et al., 2023; Chen et al., 2021b; 2023a) are enhanced and aggregated for robust representations (Lin et al., 2023; Li et al., 2024b; Tong et al., 2024). 2) High-resolution methods are proposed to overcome the fixed resolution of pre-trained vision encoders (e.g., $336\times 336$ for CLIP (Radford et al., 2021)), empowering LLMs to perceive fine-grained visual information (Liu et al., 2024a; Dong et al., 2024; Huang et al., 2024; Luo et al., 2024). 3) High-quality instruction data is essential for the VLLMs to generate accurate and well-formed responses following instructions (Laurençon et al., 2024; Li et al., 2024a; Chen et al., 2024b). In this paper, besides achieving state-of-the-art vision-language performance, we further introduce speech understanding and generating capabilities into our EMOVA.
@@
(SLLMs) empower speech interaction with LLMs. Continuous SLLMs (Wu et al., 2023; Chu et al., 2024) utilize the speech encoders (Radford et al., 2023) to extract continuous speech embeddings for LLM, which, however, only support speech understanding, relying on external TTS modules for speech generation, and therefore, hampering real-time interaction. Discrete SLLMs (Zhang et al., 2023a), instead, first discretize speech signals with speech tokenizers, followed by auto-regressive modeling. Recent works (Fang et al., 2024; Xie & Wu, 2024) further combine the continuous speech encoders with the discrete speech tokenizers for better performance. Although effective, none of the existing works explore speech style controllability in SLLMs (e.g., genders, emotions, and pitches), which is essential for real-life spoken dialogue.
@@
support visual, text, and speech capabilities with a unified architecture simultaneously. Similar to the continuous SLLMs, InternOmni (Chen et al., 2024b) and VITA (Fu et al., 2024b) connect a speech encoder with VLLMs, supporting speech understanding only. Instead, AnyGPT (Zhan et al., 2024) proposes a unified architecture to discretize and conduct auto-regressive modeling for image, text, and audio simultaneously, which, however, suffers from inevitable information loss brought by discretization, especially for the high-resolution visual inputs. In this work, we propose EMOVA, the very first unified Omni-modal Large Language Models with state-of-the-art vision-language and speech performance at the same time.
@@ 3.1 Formulation
Denote the Large Language Model (LLM) as $f(\cdot)$ and the text, visual and speech inputs as $\textbf{X}_{T}$, $\textbf{X}_{V}$ and $\textbf{X}_{S}$, respectively. $\textbf{X}_{T}$ is converted to discrete tokens $\textbf{U}_{T}$ via a text tokenizer (Gage, 1994), while the visual input $\textbf{X}_{V}$ is first encoded with a vision encoder $v(\cdot)$ as $\textbf{E}_{V}=v(\textbf{X}_{V})$, and then projected into the text embedding space with a projector $p(\cdot)$ as $\textbf{H}_{V}=p(\textbf{E}_{V})$. As for the speech input $\textbf{X}_{S}$, a Speech-to-Unit (S2U) procedure is required. Specifically, $\textbf{X}_{S}$ first passes through a speech encoder $s(\cdot)$ as $\textbf{E}_{S}=s(\textbf{X}_{S})$, which is then discretized by the quantizer $q(\cdot)$ as $\textbf{U}_{S}=q(\textbf{E}_{S})$. The LLM $f(\cdot)$ is then trained to compute the joint probability of the output text and speech units $\textbf{U}_{T}^{o},\textbf{U}_{S}^{o}$ as
$\mathbb{P}(\textbf{U}_{T}^{o},\textbf{U}_{S}^{o}|\textbf{U}_{T},\textbf{U}_{S},\textbf{H}_{V})=\prod_{i=1}^{L}\mathbb{P}(\bm{x}_{i}|\textbf{U}_{T,<i}^{o},\textbf{U}_{S,<i}^{o},\textbf{U}_{T},\textbf{U}_{S},\textbf{H}_{V}),$ (1)
where $\bm{x}_{i}\in\textbf{U}_{T}^{o}\cup\textbf{U}_{S}^{o}$ and $L=|\textbf{U}_{T}^{o}|+|\textbf{U}_{S}^{o}|$. The output response units $\textbf{U}_{S}^{o}$ are then recovered into the output speech waveform $\textbf{Y}_{S}^{o}$ via a Unit-to-Speech (U2S) decoder $d(\cdot,\cdot)$ together with an emotion style embedding $\textbf{E}_{style}^{o}$ to realize vivid emotional spoken dialogue controllability (Sec. 3.2).
@@
We adopt the LLaMA-3.1-8B (Dubey et al., 2024) as our base LLM $f(\cdot)$, due to its superior performance among publicly available checkpoints, which is equipped with a tiktoken text tokenizer and a vocabulary size of 128,256, supporting both multilingual textual inputs and outputs.
@@
We utilize InternViT-6B (Chen et al., 2024b) as our visual encoder $v(\cdot)$ with $448\times 448$ base resolution, and C-Abstractor (Cha et al., 2024) with two ResBlocks (both before and after the pooling layer) and 4×4$\times$4 × downsample rate as vision projector $p(\cdot)$. To process the high-resolution inputs, the high-resolution image-slicing (Liu et al., 2024a) is used, where visual tokens for one image are concatenation with a low-resolution thumbnail and the origin high-resolution image with separators in each line, allowing a maximum of nine tiles during training.
@@ 3.2 Speech Tokenization
Following Tao et al. (2024), we adopt the SPIRAL (Huang et al., 2022) architecture for the speech encoder $s(\cdot)$ to capture both phonetic and tonal information, which is then discretized by the quantizer $q(\cdot)$ utilizing the finite scalar quantization (FSQ)  (Mentzer et al., 2023). The size of the speech codebook is 4,096, while the sample rate is 25 tokens per second. Once discretized, the speech modality can be simply integrated into LLMs by concatenating the text vocabulary and speech codebook.
Our S2U tokenizer provides the following advantages: 1) Data efficiency: after pre-training on large-scale unlabeled speech data, it requires only a small amount of speech-text pair data for fine-tuning. 2) Bilingual: the speech codebook is shared among different languages (i.e., English and Chinese), sharing unit modeling abilities across languages. Check more training details and comparisons with other speech tokenizers (Zhang et al., 2023b) in Appendix A.1.
@@
To seamlessly align speech units with the highly semantic embedding space of LLMs, we opt for decoupling the semantic contents and acoustic styles of input speeches. Specifically, given input speechs $\textbf{X}_{S}$, both semantic embedding $\textbf{E}_{semantic}$ and style embeddings $\textbf{E}_{style}$ are extracted separately, while only the $\textbf{E}_{semantic}$ is quantified by $q(\cdot)$ to generate speech units $\textbf{U}_{S}$. By changing $\textbf{E}_{style}$ while maintaining the same $\textbf{E}_{semantic}$, we can easily control speech styles without disturbing the semantic contents of recovered speeches. Moreover, the disentanglement facilitates modality alignment among speeches and texts, as later shown in Sec. 4.1.
Unit-to-speech (U2S) detokenizer with style control.
@@
Building on VITS (Kim et al., 2021), our U2S detokenizer adopts a conditional VAE architecture (see Fig. 9). To achieve vivid style controls, we utilize the semantic-style disentanglement of our S2U tokenizer (as discussed above) and adopt a novel style embedding to control the speech styles (e.g., speaker identities, emotions, and pitches). Specifically, the LLM $f(\cdot)$ is trained to generate both the output speech units $\textbf{U}_{S}^{o}$ and a style label. The speech units $\textbf{U}_{S}^{o}$ are converted to unit embeddings $\textbf{E}_{semantic}^{o}$, while the style label is utilized to generate a unique style prototype $\textbf{E}_{style}^{o}$. Both $\textbf{E}_{semantic}^{o}$ and $\textbf{E}_{style}^{o}$ are taken as inputs to speech decoder $d(\cdot,\cdot)$ to synthesize the output speech waveform $\textbf{Y}_{S}^{o}$. See Appendix A.2 for more details.
@@
Our U2S detokenizer is pre-trained on LibriTTS (Zen et al., 2019) and AISHELL-1 (Bu et al., 2017) and subsequently fine-tuned on synthetic style-rich speech data. Specifically, due to the scarcity of real-life style-rich data, we utilize TTS tools (Du et al., 2024) to synthesize speech samples diverse in genders, pitches, and emotions. As for style prototypes, Emotion2Vec (Ma et al., 2023) is adopted to select the most representative samples with the highest confidence in conveying the desired style. Our empirical results reveal that even one representative style reference speech has been sufficient to control the speech styles flexibly and precisely.
4 Omni-modal Alignment and Instruction Tuning

### v3

# Vivid Emotions

@@ Abstract
GPT-4o, an omni-modal model that enables vocal conversations with diverse emotions and tones, marks a milestone for omni-modal foundation models. However, empowering Large Language Models to perceive and generate images, texts, and speeches end-to-end with publicly available data remains challenging for the open-source community. Existing vision-language models rely on external tools for speech processing, while speech-language models still suffer from limited or totally without vision-understanding capabilities. To address this gap, we propose the EMOVA (EMotionally Omni-present Voice Assistant), to enable Large Language Models with end-to-end speech abilities while maintaining the leading vision-language performance. With a semantic-acoustic disentangled speech tokenizer, we surprisingly notice that omni-modal alignment can further enhance vision-language and speech abilities compared with the bi-modal aligned counterparts. Moreover, a lightweight style module is introduced for the flexible speech style controls including emotions and pitches. For the first time, EMOVA achieves state-of-the-art performance on both the vision-language and speech benchmarks, and meanwhile, supporting omni-modal spoken dialogue with vivid emotions.
@@ 1 Introduction
OpenAI GPT-4o [70], a novel milestone for the omni-modal foundation models, has rekindled people’s attention on intelligent assistants that can see (i.e., perceiving fine-grained visual inputs), hear (i.e., understanding vocal instructions) and speak (i.e., generating vocal responses) simultaneously. Existing Multi-modal Large Language Models (MLLMs) mostly focus on two modalities only, either vision-language [1, 40] or speech-language [11, 87], demonstrating severe demands for omni-modal models with visual, language and speech abilities. How to empower Large Language Models (LLMs) to effectively process omni-modal data in an end-to-end manner remains an open question.
Figure 1: EMOVA is the very first omni-modal LLM with state-of-the-art performance on both vision-language and speech benchmarks simultaneously. See detailed results in Table 2.
@@
Existing omni-modal LLMs [10, 19] generally build on Vision LLMs and integrate the speech modality by adopting a speech encoder like Whisper [74], which extracts continuous speech features, similar to how images are processed, to enable speech understanding. These models, however, still rely on external Text-to-Speech (TTS) tools for generating speech responses, limiting their ability to support real-time interactions. AnyGPT [92], on the contrary, opts for a fully discretization way, which first discretizes all data modalities (i.e., images, texts, and speeches), followed by omni-modal auto-regressive modeling. This enables AnyGPT to handle multiple modalities with a unified end-to-end framework, facilitating real-time interactions with the help of streaming decoding. However, the discrete vision tokenizer used by AnyGPT struggles to capture visual details, especially for high-resolution images, falling far behind its continuous counterparts on vision-language benchmarks. Furthermore, none of the existing works explore the speech style controls (e.g., emotions and pitches) with LLMs. Thus, our question arises: How to build an end-to-end omni-modal LLM enabling emotional spoken dialogue while maintaining state-of-the-art vision-language performance?
@@
In this paper, we propose EMOVA (EMotionally Omni-present Voice Assistant), a novel end-to-end omni-modal LLM with state-of-the-art vision-language and speech capabilities while supporting emotional spoken dialogue. Fig. 2 shows an overview of the model framework. A continuous vision encoder captures fine-grained visual details, while a discrete speech tokenizer and detokenizer empower end-to-end speech understanding and generation. Specifically, the speech-to-unit (S2U) tokenizer tokenizes the input speech waveforms to discrete speech units, while the unit-to-speech (U2S) detokenizer reconstructs the speech waveforms from LLM’s output speech units. To seamlessly integrate speech modality with LLMs, we meticulously design a semantic- acoustic disentangled speech tokenizer to decouple the semantic contents and acoustic styles of input speeches [77], where 1) semantic content (i.e., what it says) captures the semantic meanings of the input speeches, which is finally discretized and aligned with LLMs, while 2) acoustic style (i.e., how it says) captures the diverse speech styles (e.g., emotions and pitches). Utilizing the semantic-acoustic disentanglement of our speech tokenizer, we further introduce a lightweight style module to support spoken dialogue with vivid emotions and pitches. As in Sec. 4.1, this disentanglement design better facilitates the modality alignment among texts and speeches while maintaining flexibility for diverse speech style controllability and personalization.
@@
With EMOVA’s end-to-end omni-modal framework, we empirically demonstrate publicly available bi-modal image-text and speech-text data are sufficient for the omni-modal alignment with the text modality as a bridge, eliminating the need for omni-modal data (i.e., image-text-speech), which is usually scarce. Surprisingly, we find that the omni-modal alignment can further improve both the vision-language and speech capabilities via joint optimization, even when compared with their bi-modal aligned counterparts. Ultimately, only a small amount of omni-modality samples are required to teach the model to respond in the desired format. For the first time, EMOVA obtains state-of-the-art results on both vision-language and speech benchmarks (see Table 2). The main contributions of this work contain three parts:
@@
    We propose EMOVA, a novel end-to-end omni-modal LLM that can see, hear and speak. A continuous vision encoder with a semantic-acoustic disentangled speech tokenizer is adopted for seamless omni-modal alignment and diverse speech style controllability.
@@
    We introduce an efficient text-centric omni-modal alignment which can further enhance the vision-language and speech abilities, surpassing their bi-modal aligned counterparts (i.e., image-text only and speech-text only).
@@
    For the first time, EMOVA obtains state-of-the-art comparable results on both the vision-language and speech benchmarks simultaneously and further supports flexible spoken dialogues with vivid emotions.
@@
Understand Gen. Emotion
Vision Large Language Models
LLaVA [49]
Intern-VL [10]
Speech Large Language Models
Qwen-Audio [11]
Mini-Omni [87]
LLaMA-Omni [17]
Omni-modal Large Language Models
VITA [19, 20]
Ola [55]
Any-GPT [92]
Baichuan-Omni [43]
EMOVA (ours)
Table 1: Comparison of Multi-modal Large Language Models. EMOVA is the very first Omni-modal LLM capable of emotional spoken dialogue with state-of-the-art vision-language and speech capabilities simultaneously. “Gen.” stands for Generation.
@@ 2 Related Work
(VLLMs) integrate the vision modality into LLMs [81, 7], enabling the advanced understanding and reasoning over visual instructions [49, 1, 26, 27]. Recent VLLM works can be categorized into three directions, 1) Vision encoders [71, 5, 6] are enhanced and aggregated for robust representations [46, 45, 80]. 2) High-resolution methods are proposed to overcome the fixed resolution of pre-trained vision encoders (e.g., $336\times 336$ for CLIP [73]), enabling LLMs to perceive fine-grained visual information [48, 14, 31, 58]. 3) High-quality instruction data is essential for VLLMs to generate accurate and well-formed responses [36, 40, 10]. Besides achieving state-of-the-art vision-language performance, we further introduce speech understanding and generating abilities to EMOVA.
Figure 2: Model architecture of EMOVA. The vision encoder extracts continuous visual features, which are projected into the textual embedding space as visual tokens, while the input speech is encoded and quantized into discrete speech units. Given the omni-modal inputs, EMOVA can generate both textual and speech responses with vivid emotional controls. Check Sec. 3 for more architectural details.
@@
(SLLMs) empower the speech interaction with LLMs. Continuous SLLMs [84, 11] adopt the speech encoders [74] to extract continuous speech embeddings for LLM, which, however, only support speech understanding, relying on external TTS modules for speech generation, and therefore, hampering real-time interaction. Discrete SLLMs [94], instead, first discretize speech signals with speech tokenizers, followed by auto-regressive modeling. Recent works [17, 87] further combine the continuous speech encoders with discrete speech tokenizers for better results. Although effective, none of the existing works explore speech style controllability in SLLMs (e.g., emotions and pitches), which is essential for spoken dialogue.
@@
support visual, text, and speech abilities with a unified architecture. Similar to continuous SLLMs, InternOmni [10] and VITA [19] connect a speech encoder with VLLMs, supporting speech understanding only. Instead, AnyGPT [92] proposes a unified architecture to discretize and conduct auto-regressive modeling for image, text, and audio simultaneously, which, however, suffers from inevitable information loss brought by discretization, especially for the high-resolution visual inputs. Our EMOVA is the very first unified Omni-modal LLM with state-of-the-art vision-language and speech performance at the same time.
@@ 3.1 Formulation
Denote LLM as $f(\cdot)$ and text, visual and speech inputs as $\textbf{X}_{T}$, $\textbf{X}_{V}$ and $\textbf{X}_{S}$, respectively. $\textbf{X}_{T}$ is converted to discrete tokens $\textbf{U}_{T}$ via a text tokenizer [21], while the $\textbf{X}_{V}$ is first encoded with a vision encoder $v(\cdot)$ as $\textbf{E}_{V}=v(\textbf{X}_{V})$, and then projected into the textual embedding space with a projector $p(\cdot)$ as $\textbf{H}_{V}=p(\textbf{E}_{V})$. As for the speech input $\textbf{X}_{S}$, a Speech-to-Unit (S2U) procedure is required. Specifically, $\textbf{X}_{S}$ first goes through a speech encoder $s(\cdot)$ as $\textbf{E}_{S}=s(\textbf{X}_{S})$, which is then discretized by the quantizer $q(\cdot)$ as $\textbf{U}_{S}=q(\textbf{E}_{S})$. The LLM $f(\cdot)$ is then trained to compute the joint probability of the output texts $\textbf{U}_{T}^{o}$ and speech units $\textbf{U}_{S}^{o}$ as
$\mathbb{P}(\textbf{U}_{T}^{o},\textbf{U}_{S}^{o}|\textbf{U}_{omni})=\prod_{i=1}^{L}\mathbb{P}(\boldsymbol{x}_{i}|\textbf{U}_{T,<i}^{o},\textbf{U}_{S,<i}^{o},\textbf{U}_{omni}),$ (1)
where $\boldsymbol{x}_{i}\in\textbf{U}_{T}^{o}\cup\textbf{U}_{S}^{o}$, $L=|\textbf{U}_{T}^{o}|+|\textbf{U}_{S}^{o}|$ and $\textbf{U}_{omni}=\textbf{U}_{T}\cup\textbf{U}_{S}\cup\textbf{H}_{V}$, which stands for the omni-modal inputs. The output response units $\textbf{U}_{S}^{o}$ are then recovered into the output speech waveform $\textbf{Y}_{S}^{o}$ via a Unit-to-Speech (U2S) decoder $d(\cdot,\cdot)$ with an emotion style embedding $\textbf{E}_{style}^{o}$ to realize the vivid emotional spoken dialogue controllability (Sec. 3.2).
@@
We utilize the Qwen-2.5 [79] model families as the base LLMs of EMOVA with three configurations (i.e., 3B, 7B, and 72B) for usage under different budgets.
@@
We use the QwenViT [78] as the visual encoder $v(\cdot)$ with an MLP vision projector $p(\cdot)$ with a 4×4$\times$4 × downsample rate for all variants of EMOVA.
[…]
@@ 3.2 Speech Tokenization
Following [77], we use the SPIRAL [32] architecture for the speech encoder $s(\cdot)$ to capture both phonetic and tonal information, which is then discretized by the quantizer $q(\cdot)$ with finite scalar quantization (FSQ) [64]. The size of the speech codebook is 4,096, while the sample rate is 25 tokens per second. Once discretized, the speech modality can be integrated into LLMs by concatenating the text vocabulary and speech codebook.
Our S2U tokenizer provides the following advantages: 1) Data efficiency: after pre-training on large-scale unlabeled speech data, it requires only a small amount of speech-text pair data for easy adaptation. 2) Bilingual: the speech codebook is shared among languages (i.e., English and Chinese), sharing unit modeling capabilities across languages. Check more training details and comparisons in Appendix A.1.
@@
To align the speech units seamlessly with the highly semantic embedding space of LLMs, we opt for decoupling the semantic contents and acoustic styles of input speeches. Given input speechs $\textbf{X}_{S}$, both semantic embedding $\textbf{E}_{semantic}$ and style embeddings $\textbf{E}_{style}$ are extracted separately as
$\{\textbf{E}_{semantic},\textbf{E}_{style}\}=\textbf{E}_{S}=s(\textbf{X}_{S}).$ (2)
Only $\textbf{E}_{semantic}$ is quantified by $q(\cdot)$ to generate the speech units $\textbf{U}_{S}$. By using different $\textbf{E}_{style}$ while maintaining the same $\textbf{E}_{semantic}$, we can easily control the recovered speech styles without disturbing the semantic contents of recovered speeches. Moreover, the disentanglement facilitates modality alignment among speeches and texts, as later in Sec. 4.1.
Unit-to-speech (U2S) detokenizer with style controls.
@@
Building on VITS [35], our U2S detokenizer adopts a conditional VAE architecture (cf., Fig. 7). To achieve flexible style controls, we utilize the semantic-acoustic disentanglement of our S2U tokenizer (as discussed above) and adopt a novel style embedding to control the speech styles (e.g., genders, emotions, and pitches). Specifically, the LLM $f(\cdot)$ is trained to generate both the output speech units $\textbf{U}_{S}^{o}$ and a style label. The speech units $\textbf{U}_{S}^{o}$ are converted to the unit embeddings $\textbf{E}_{semantic}^{o}$, while the style label is utilized to generate a unique style prototype $\textbf{E}_{style}^{o}$. Both $\textbf{E}_{semantic}^{o}$ and $\textbf{E}_{style}^{o}$ are taken as inputs to speech decoder $d(\cdot,\cdot)$ to synthesize output speeches $\textbf{Y}_{S}^{o}=d(\textbf{E}_{semantic}^{o},,\textbf{E}_{style}^{o})$.
@@
Our U2S detokenizer is pre-trained on LibriTTS [91] and AISHELL-1 [2] and subsequently fine-tuned on synthetic style-rich speech data. Due to the scarcity of real-life style-rich data, we utilize TTS tools [16] to synthesize the speech samples diverse in genders, pitches, and emotions. As for the style prototypes, Emotion2Vec [59] is adopted to select the most representative samples with the highest confidence in conveying the desired style. Our empirical results reveal that even one representative style reference speech has been sufficient to control the speech styles flexibly and precisely. Check Appendix A.2 for more details.
Figure 4: Demonstration of EMOVA omni-modal instruction tuning. 1) To support emotional spoken dialogues, EMOVA is trained to explicitly select speech style labels with output speech units. 2) For ease of parsing, the data elements are organized in the JSON format.
4 Training Omni-modal LLMs